A slightly bruised apple will rot extensively in about 2.0 days at room temperature (26.C). If it is kept in the refrigerator at 1.5 C, the same extent of rotting takes about 17 days. What is the activation energy from the rotting reaction?          ......................     J/mol             fyi( 5.9 J/mol was marked as wrong)
Answer #1

Using the equation,

ln(k1/k2) = Ea/R(1/T2-1/T1)

where,

Ea = activation energy

k1 = 2

k2 = 17

T1 = 299 K

T2 = 274.5 K

R = 8.314 J/K.mol

Feed values,

ln(2/17) = Ea/8.314[1/299-1.274.5]

Ea = 59603.43 J/mol